OpenStudy (anonymous):

Hi everyone! Can anyone confirm my answer of y=1/4x^2 + 2x +c for the DE dy/dx=2+root(y-2x+3) and use a u-substitution of u=y-2x+3...Thanks! :o)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (perl):

is this the problem? $$ \Large \frac{dy}{dx}= 2 + \sqrt{y-2x + 3} $$

OpenStudy (anonymous):

@perl yes...sorry, got tied up for awhile

OpenStudy (michele_laino):

I got this general integral: \[y\left( x \right) = \frac{{{{\left( {x + c} \right)}^2}}}{2} + 2x - 3\]

OpenStudy (anonymous):

hmmm...you don't have your work handy do you? The reason I ask is because I worked this twice and got the same answer... I very well could be wrong but I'm not 100% sure your answer is correct either

OpenStudy (michele_laino):

oops...: \[y\left( x \right) = {\left( {\frac{{x + c}}{2}} \right)^2} + 2x - 3\]

OpenStudy (michele_laino):

using your substitution, I get: \[u' = \sqrt u \]

OpenStudy (michele_laino):

so, integrating once, I can write: \[2\sqrt u = x + c\]

OpenStudy (anonymous):

yikes...I had this problem on my whiteboard and erased it! :o/ I will work it again on paper and post it if you have time to compare it with your work?

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

thanks! :o) brb

OpenStudy (anonymous):

One of the way to confirm: if y = 1/4 x^2 +2x +c is solution , then y' =........ consider given y'= 2+root(y -2x+3), replace y above into it, if it is = y' we calculate above, then we are correct.